OpenStudy (anonymous):

the correct simplification of the expression (2ab5)3

OpenStudy (anonymous):

its being squared

jimthompson5910 (jim_thompson5910):

\[(2ab^5)^3=(2ab^5)(2ab^5)(2ab^5)=(2*2*2)*(ab^5*ab^5*ab^5)=8a^3b^{15}\]

OpenStudy (aroub):

8a^3b^15